# SynWorld: Virtual Scenario Synthesis for Agentic Action Knowledge Refinement

### v1

@@ Abstract
In the interaction between agents and their environments, agents expand their capabilities by planning and executing actions. However, LLM-based agents face substantial challenges when deployed in novel environments or required to navigate unconventional action spaces. To empower agents to autonomously explore environments, optimize workflows, and enhance their understanding of actions, we propose SynWorld, a framework that allows agents to synthesize possible scenarios with multi-step action invocation within the action space and perform Monte Carlo Tree Search (MCTS) exploration to effectively refine their action knowledge in the current environment. Our experiments demonstrate that SynWorld is an effective and general approach to learning action knowledge in new environments. [^1] Code is available at https://github.com/zjunlp/SynWorld.
@@ A.2 Knowledge-Augmented Agents
LLMs, as agents interacting with specific environments, often need to provide action signals to these environments Zhou et al. (2023, 2024); Durante et al. (2024). These action signals can be either restricted or open actions related to the environment Wang et al. (2024); Li et al. (2024); Hu et al. (2024). For instance, they might involve specific movements in embodied task scenarios or the use of various tools like search or OCR in tool invocations. By incorporating these actions, on one hand, the Agent gains the ability to interact with the environment, allowing LLMs to transcend mere textual output Shridhar et al. (2021); Wang et al. (2022). On the other hand, these external actions endow the Agent with a capability similar to humans using tools, compensating for the inherent limitations of LLMs, such as search tools that can alleviate issues of knowledge hallucination or obsolescence in LLMs Singh et al. (2024); Zhao et al. (2024b); Shen et al. (2024).

### v3

@@ Abstract
In the interaction between agents and their environments, agents expand their capabilities by planning and executing actions. However, LLM-based agents face substantial challenges when deployed in novel environments or required to navigate unconventional action spaces. To empower agents to autonomously explore environments, optimize workflows, and enhance their understanding of actions, we propose SynWorld, a framework that allows agents to synthesize possible scenarios with multi-step action invocation within the action space and perform Monte Carlo Tree Search (MCTS) exploration to effectively refine their action knowledge in the current environment. Our experiments demonstrate that SynWorld is an effective and general approach to learning action knowledge in new environments [^1] Code is available at https://github.com/zjunlp/SynWorld..
@@ 4.1 Experiment Setup
Qwen2-72B-Instruct ReAct 49.43 55.00 50.21
Self-Refine 54.33 65.00 52.59
EasyTool 52.97 58.00 54.94
DRAFT 56.43 69.00 57.57
Ours 58.52 73.00 58.70
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), CIPSC-SMP-Zhipu Large Model Cross-Disciplinary Fund, Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ A.2 Knowledge-Augmented Agents
LLMs, as agents interacting with specific environments, often need to provide action signals to these environments Zhou et al. (2023, 2024); Durante et al. (2024). These action signals can be either restricted or open actions related to the environment Wang et al. (2024); Li et al. (2024); Hu et al. (2024); Zhang et al. (2024). For instance, they might involve specific movements in embodied task scenarios or the use of various tools like search or OCR in tool invocations. By incorporating these actions, on one hand, the Agent gains the ability to interact with the environment, allowing LLMs to transcend mere textual output Shridhar et al. (2021); Wang et al. (2022); Zhao et al. (2024b). On the other hand, these external actions endow the Agent with a capability similar to humans using tools, compensating for the inherent limitations of LLMs, such as search tools that can alleviate issues of knowledge hallucination or obsolescence in LLMs Singh et al. (2024); Zhao et al. (2024c); Shen et al. (2024); Chen (2023).